# In Situ AI Prototyping: Infusing Multimodal Prompts into Mobile Settings with MobileMaker

### v1

@@ Abstract
Recent advances in multimodal large language models (LLMs) have lowered the barriers to rapidly prototyping AI-powered features via prompting, especially for mobile-intended use cases. Despite the value of situated user feedback, the process of soliciting early, mobile-situated user feedback on AI prototypes remains challenging. The broad scope and flexibility of LLMs means that, for a given use-case-specific prototype, there is a crucial need to understand the wide range of in-the-wild input likely to be provided by the user, as well as their in-context expectations of the AI’s behavior. To explore the concept of in situ AI prototyping and testing, we created MobileMaker: an AI prototyping tool that enables designers to rapidly create mobile AI prototypes that can be tested on-device, and enables testers to make on-device, in-the-field revisions of the prototype through natural language. In an exploratory study with 16 users, we explored how user feedback on prototypes created with MobileMaker compares to that of existing prototyping tools (e.g., Figma, prompt editors). We found that MobileMaker prototypes enabled more serendipitous discovery of: model input edge cases, discrepancies between AI’s and user’s in-context interpretation of the task, and contextual signals missed by the AI. Furthermore, we learned that while the ability to make in-the-wild revisions led users to feel more fulfilled as active participants in the design process, it might also constrain their feedback to the subset of changes perceived as more actionable or implementable by the prototyping tool.
@@ I Introduction
Recent advances in multimodal large language models (LLMs) are lowering the barriers to rapidly prototyping novel AI-powered features and products via prompt programming [1, 2, 3, 4]. Compared to the traditional, more time-consuming AI development cycle of data collection, model training, and integration into UIs [5, 6], it is now much easier to prototype AI through “prompt-based prototyping” [7, 8, 9]: crafting LLM prompts in minutes and embedding them into new AI prototypes. Simultaneously, the multimodality of recent LLMs (e.g., GPT4 [10], Gemini [11]) has made it possible to process rich inputs (e.g., text, images, video), unlocking a wide range of use cases suitable for mobile or “on the go” settings, from diagnosing plant diseases (from a photo and text list of symptoms) to generating recipes (based on a picture of ingredients and set of dietary restrictions).
Despite this rapid pace of LLM advancement and AI prototyping, the process of soliciting early user feedback on AI prototypes remains challenging. Prior work on “embodied” user feedback [12, 13, 14] has shown that real-world, in-the-wild settings lead to more authentic user feedback than lab settings because prototypes are experienced within the final application’s intended environment. The specific idiosyncrasies and flexibility of LLMs make it even more crucial to situate LLM-powered AI prototypes and user feedback within authentic environments. Whereas traditional machine learning models were scoped to specific, narrow applications, with inputs and outputs relatively known a priori [15, 16, 17], LLMs and foundational models can be applied to a wide range of tasks (e.g., code completion, medical help, tutoring) [18]; receive a wide range of “natural” real-world inputs (e.g., photography, icons, diagrams, text, etc.) [19]; and potentially be applied in a wide range of settings (e.g., home, commute, work, coffee shop). Given this broad scope, for a given use case-specific prototype, there is the need to understand the types of input likely to be provided by the user, their in-context expectations of the AI’s behavior, and validation of the overall concept.
@@
LLMs also open up new opportunities for in situ prototyping and testing. First, because prompt programming has significantly reduced the ML expertise required to prototype AI, there is an opportunity for designers to more easily bring LLM-powered prototypes to life within mobile user interfaces (UIs) for in in-the-wild experiences. Secondly, much as paper prototyping enables users and designers alike to alter designs on-the-fly, LLM-powered prototypes could enable testers to proactively revise and improve prototypes during testing, at the very moments they encounter an issue or get an idea for an improvement to the design.
@@
To explore these opportunities for in situ AI prototyping and testing, we created MobileMaker: an AI prototyping tool that 1) enables designers to rapidly create a mobile AI prototype that can be tested on device, and 2) enables testers to make on-device, in-the-field revisions of the prototype’s design through natural language. As such, MobileMaker allows users to provide not only feedback, but also alternative, functional prototypes that better match their specific needs. In a user study with 16 participants, we explored how user feedback on AI prototypes created on MobileMaker compares to that of existing AI prototyping tools (e.g., Figma, LLM prompt editors). In a follow-up reflection activity, two professional designers envisioned how the user feedback and revisions collected could influence the iterative design process.
@@
We found that MobileMaker prototypes enabled users to authentically experience the envisioned artifact, which allowed them to 1) evaluate if the overall concept was mobile appropriate, 2) experiment with serendipitous edge cases via unconventional, in-the-wild inputs, 3) discover discrepancies between their interpretation of the task and the AI’s interpretation, and 4) more critically evaluate the AI’s output using the richer contextual cues in their surroundings. However, testers couldn’t always quickly place themselves in representative contexts due to scheduling and location constraints, which might have limited their ability to provide grounded feedback.
We also found that the ability to make in-the-wild prototype revisions led users to feel more fulfilled as active participants in the design process and more critically assess their own feedback. Furthermore, though designers were initially hesitant to let users “short-circuit” the design process, they found the revised prototypes to be a more powerful and convincing form of feedback, especially because they felt traditional user feedback could be more easily overlooked or dismissed. Finally, while on-device revisions could potentially shorten the test-feedback design loop, user considerations around what revisions were actionable may have constrained users’ higher-level feedback.
In sum, this paper makes the following contributions:
@@
  • Design goals for LLM-powered mobile AI prototyping platforms, based on formative studies and iterative feedback from designers
@@
  • MobileMaker, a prototyping tool that enables designers to create, test with users, and revise AI prototypes in natural language while on mobile
@@
  • An exploratory user study that uncovers how in situ prototypes (created using MobileMaker) led to qualitatively different user feedback, compared to that of existing AI prototypes (e.g., UI mockups, LLM prompt editors)
@@
  • A follow-up reflection exercise with two designers, showing how designers can make use of fully functional revised prototypes created by users in the wild
@@
  • A discussion of implications for future design and prototyping platforms, advocating for systems that prioritize quick iterations and authentic user feedback for AI experienced in-the-wild
@@ II-B In Situ User Testing
After developing a prototype, it is crucial to perform user testing and gather feedback [26, 42, 12]. Traditional UI prototyping tools like Figma [43] allow users to interact with click-through prototypes [43, 44, 45], but these often include only pre-defined input and output examples [22, 46], and the quality and relevance of these examples may not fully represent the diverse contexts and challenges that end-users may face in the real-world [47]. To circumvent the technical demands of building functional AI prototypes while still engaging end-users, designers often conduct user testing with Wizard of Oz setups [48], where a human simulates AI responses [49, 50, 51]. Although helpful, this often fails to accurately simulate the unique errors and behaviors of actual AI systems [52, 53].
To address these challenges, in this work, we focus on supporting “in situ” user testing of AI prototypes, i.e., enabling interactions with MobileMaker prototypes in their intended real-world environments [14, 13, 54, 55]. Specifically, MobileMaker allows users to capture inputs from their immediate surroundings using their mobile devices and view AI model outputs in that same context. We found that this approach not only facilitated the discovery of serendipitous edge cases from real-world inputs, but also enabled users to evaluate the AI’s performance more critically, leveraging the rich contextual cues around them.
@@ II-C Iterative Design Based on User Feedback
In HCI, iterative design is recognized as a critical method that emphasizes the repeated cycle of designing, testing, and refining products based on user feedback [56, 57, 58], which is crucial in helping designers identify user pain points and areas for improvements [57, 59]. The length of each iteration in the design process can vary significantly–ranging from a few days in rapid prototyping scenarios to several weeks or months in more sophisticated systems–based on factors like the iteration’s specific objectives, the level of fidelity of the prototype involved, and the available resources [54, 60, 61, 62]. In this work, we explore the opportunity to accelerate these iteration cycles even further by having MobileMaker instantly implement users’ feedback on the spot, thereby enabling them to immediately interact with a revised version of the prototype. This allowed our user study participants to more critically reflect on their own feedback and increased their perceived engagement in the design process. Moreover, users can submit these updated prototypes as “revisions” back to designers, providing a more direct means of conveying user needs in specific contexts.
@@ III Formative Studies & Design Goals
Over several months, we conducted a series of needfinding interviews with four professional UX designers at a large technology company. Three out of four had experience in designing and prototyping AI and LLM-powered features, and one also had extensive experience creating high-fidelity web-based prototypes powered by LLMs. We inquired about the interviewees’ current workflows of designing, prototyping, and testing AI-powered features, as well as challenges they encountered. In each session, participants were asked to envision and prototype an AI-powered feature using an earlier version of MobileMaker. We monitored their usage, collecting insights and suggestions for improvement.
@@
Based on the data from these formative interviews as well as prior work, we identified three design goals for MobileMaker:
@@
  • D1: Rapid and frictionless authoring of medium-fidelity, mobile AI prototypes. Designers typically prototype AI functionality and UI separately and with different tools. Integrating the AI and UI into a working prototype usually requires significant engineering resources and could often “take days if not weeks” with considerable coordination and communication overhead. Most designers did not attempt to put early-stage prototypes on mobile phones due to the significant friction involved (e.g., security, on-device configuration). While using early versions of MobileMaker, designers were excited about the possibility of being able to quickly wrap LLM-powered prompts “in an app shell.” They noted that this would substantially accelerate prototyping velocity and reduce their reliance on engineers. Some were also particularly interested in the ability to transform ideas into implementations using natural language. The designers clarified that they did not envision this platform to be for creating high-fidelity, pixel-perfect prototypes or fully-fleshed mobile applications. Instead, they wanted prototypes that could convey the concept and were functional enough to help uncover target user intents, potential inputs, and reflections on the model’s outputs.
  • D2: Lower the barriers to on-device user experience and feedback on new AI-powered ideas. Designers currently solicit AI prototype feedback through UI mockups like Figma or showing example model inputs and outputs in prompt editors. Designers emphasized that MobileMaker-style prototypes could be more approachable to early users and testers (compared to prompt editors) and more dynamically realistic in its approximation of the AI (compared to UI mockups), enabling users to experience a better approximation of the envisioned artifact while out and about on their phones.
@@
  • D3: Expedited design iteration loop. Traditionally, refining a mobile prototype after user testing is a lengthy and resource-intensive process: designers must make revisions, collaborate with engineers to update the prototype, and schedule new testers, all of which can take days or weeks. To streamline this, designers envisioned a future where the feedback loop is drastically shortened. Designers were excited about the possibility of having testers offer feedback on the initial prototype, then having a near-instant mechanism to produce an improved version of the prototype for another round of testing, without needing to wait days or weeks for an engineer to make updates.
@@ IV MobileMaker
Based on the above design rationales, we describe the system and user interface design of MobileMaker. Our formative studies revealed that designers sought to make prototypes capable of unveiling user intentions, inputs, and reflections on model outputs without being overly complex. Thus, our objective was not to develop a platform for authoring fully-functional mobile applications; instead, we focused on implementing a minimal set of features essential for exploring the human-AI interaction of the prototypes.
@@ IV-A Creating Prototypes
In MobileMaker, users can build prototypes using three types of configurable UI widgets (Fig. 2):
  • Input Widgets represent different modalities of input that users can provide, including text, dropdown menus (with a list of predefined text values), live-camera photos (captured live from the front or back camera feed on the mobile device), and file uploads (e.g., photos from an on-device album) (Fig. 2A)
@@
  • Output Widgets are used to display the output from generative models, i.e., LLMs, multimodal LLMs, and image generation models  [^1] MobileMaker uses the Vertex AI Gemini and Imagen APIs [63, 64]. (Fig. 2C)
@@
Each widget type offers customizable visual elements, such as text input placeholders (e.g., “Enter a color”), widget labels (e.g., “Upload a photo of your room”), and the choice to display model outputs as plain text or in a visual component (e.g., carousel card). Users can also configure various global parameters, include the prototype’s name, description, font style, and layout style, i.e., whether camera controls should be laid out in the list of input controls or as a full screen layout behind all other controls.
@@ IV-B Prototyping with Natural Language
From early feedback, we discovered that designers appreciate the capability to quickly convert ideas into functional mobile prototypes (D1)–despiteMobileMaker’s straightforward interface, some designers found translating high-level concepts to specific low-level widgets cognitively demanding, and could hinder “spur of the moment” ideas. To reduce this barrier, MobileMaker additionally supports prototyping through two natural language (NL) features:
IV-B1 Natural Language Prototype Creation
In order to allows users to quickly bootstrap a prototype from NL, they can simply provide a brief sketch of their prototype idea (e.g., “a feature that generates a music playlist from a photo the user takes”). The system then uses a few-shot LLM prompt (see Appendix A1) to instantly generate a prototype. Each example in the few-shot prompt consists of an input NL request and output JSON specification, which encompasses configurations for input, action, and output widgets and as detailed previously. Critically, output widget prompts are dynamically generated to reference the corresponding generated input widgets, reducing the need for further manual configuration.
IV-B2 Natural Language Prototype Revision
To allow users to rapidly update and revise prototypes (e.g., in response to in-the-wild inspirations), MobileMaker also supports NL revisions (Fig. 2E). To achieve this, they can simply provide a brief description of the desired changes, such as “Add a dropdown for music genre.” The system then categorizes the request as either an update to an existing output widget prompt (see appendix A2) or an adjustment to the prototype structure (e.g., adding or removing widgets). For the former, the system executes a prompt-revision meta-prompt (see appendix A3) designed to modify an existing prompt based on a user request. For the latter, the system combines the user’s request and the current prototype JSON in a few-shot prompt (similar to the one for NL prototype creation; see appendix A5), which outputs a revised JSON specification for the entire prototype.
When a user requests an NL revision, they receive a summary outlining the changes between the original prototype and the revised versions. They can toggle between these two versions and decide whether to accept the revision. This feature makes it easy for users to quickly compare and confirm if their requested revision has indeed been implemented, enhancing their ability to confidently and efficiently iterate on their design.
Together, these NL features enable rapid development and require minimal manual configuration, making MobileMaker’s medium-fi prototyping process accessible and approachable to not only designers but also users with limited prototyping or prompting experience.

### v2

@@ Abstract
Recent advances in multimodal large language models (LLMs) have made it easier to rapidly prototype AI-powered features, especially for mobile use cases. However, gathering early, mobile-situated user feedback on these AI prototypes remains challenging. The broad scope and flexibility of LLMs means that, for a given use-case-specific prototype, there is a crucial need to understand the wide range of in-the-wild input users are likely to provide and their in-context expectations for the AI’s behavior. To explore the concept of in situ AI prototyping and testing, we created MobileMaker: a platform that enables designers to rapidly create and test mobile AI prototypes directly on devices. This tool also enables testers to make on-device, in-the-field revisions of prototypes using natural language. In an exploratory study with 16 participants, we explored how user feedback on prototypes created with MobileMaker compares to that of existing prototyping tools (e.g., Figma, prompt editors). Our findings suggest that MobileMaker prototypes enabled more serendipitous discovery of: model input edge cases, discrepancies between AI’s and user’s in-context interpretation of the task, and contextual signals missed by the AI. Furthermore, we learned that while the ability to make in-the-wild revisions led users to feel more fulfilled as active participants in the design process, it might also constrain their feedback to the subset of changes perceived as more actionable or implementable by the prototyping tool.
@@ I Introduction
Recent advances in LLMs have lowered the barriers to rapidly prototyping novel AI-powered features and products via prompt programming [1, 2, 3, 4]. Compared to the traditional, more time-consuming AI development cycle of data collection, model training, and integration into UIs [5, 6], it is now much easier to prototype AI through “prompt-based prototyping” [7, 8, 9]: crafting LLM prompts in minutes and embedding them into new AI prototypes. Simultaneously, the multimodality of recent LLMs (e.g., Gemini [10], GPT4 [11]) has made it possible to process rich inputs (e.g., text, images, video), unlocking a wide range of use cases suitable for mobile or “on the go” settings, from diagnosing plant diseases (from a photo and text list of symptoms) to generating recipes (based on a picture of ingredients and set of dietary restrictions).
Despite rapid advancement of LLMs and AI prototyping, gathering early tester feedback on AI prototypes remains challenging, especially in mobile environments. Previous research on “embodied” user feedback [12, 13, 14] suggests that real-world, in-the-wild settings yield more authentic tester feedback than controlled lab settings, as prototypes are used within their intended context. This is especially crucial for LLM-powered AI prototypes due to LLMs’ specific idiosyncrasies and flexibility. Unlike traditional machine learning models, which are typically scoped to specific, narrow applications with inputs and outputs relatively known a priori [15, 16, 17], LLMs can be applied to a broad spectrum of tasks (e.g., code completion, medical help, tutoring), receive diverse “natural” inputs (e.g., photography, icons, diagrams, text, audio), and function in various real-world settings (e.g., home, commute, work, coffee shop). Given this broad scope, for a given use case-specific prototype, there is the need to understand the types of input testers are likely to provide, their context-specific expectations of the AI’s behavior, and validation of the overall prototype concept.
@@
LLMs also open up new opportunities for in situ prototyping and testing. First, because prompt programming has significantly reduced the ML expertise required to prototype AI, there is an opportunity for designers to more easily bring LLM-powered prototypes to life within mobile user interfaces (UIs) for in-the-wild experiences. Secondly, much as paper prototyping enables designers and testers [^1] In this paper, we define “designers” as those who create prototypes, and “testers” as individuals who use these prototypes and provide user feedback. to alter designs on-the-fly, LLM-powered prototypes could enable testers to proactively revise and improve prototypes during testing, at the very moments they encounter an issue or get an idea for an improvement to the design.
@@
To explore these opportunities for in situ AI prototyping and testing, we created MobileMaker: an AI prototyping tool that 1) enables designers to rapidly create a mobile AI prototype that can be tested on-device, and 2) enables testers to revise the prototype’s design in-the-wild simply through natural language (NL). Specifically, MobileMaker allows testers to provide not only feedback, but also alternative, functional prototypes that better meet their specific needs and can be instantly re-tested. We conducted an exploratory study to examine how tester feedback using MobileMaker (on functional LLM-powered prototypes) compares to and complements traditional user-feedback mechanisms, and to assess the impact of in-the-wild revisions on the testing experience. In a subsequent reflection activity, two professional designers discussed how the tester-generated revisions could influence the iterative design process.
@@
We found that MobileMaker prototypes enabled testers to authentically experience the envisioned artifact, which allowed them to 1) evaluate if the overall concept was mobile appropriate, 2) experiment with serendipitous edge cases via unconventional, in-the-wild inputs, 3) discover discrepancies between their interpretation of the task and the AI’s interpretation, and 4) more critically evaluate the AI’s output using the richer contextual cues in their surroundings. We also discovered that the ability to make on-the-spot prototype revisions led testers to feel more engaged as active participants in the design process and more critically assess their own feedback. Furthermore, though designers were initially hesitant to let testers “short-circuit” the design process, they later recognized revised prototypes as a more powerful and convincing form of feedback, especially because they felt traditional user feedback could be more easily overlooked or dismissed. In sum, this paper makes the following contributions:
@@
  • Design goals for LLM-powered mobile AI prototyping platforms, based on formative studies and iterative feedback from designers,
@@
  • MobileMaker, a prototyping tool that enables designers to create, test, and revise AI prototypes in NL while on mobile,
@@
  • an exploratory study showing how in situ prototypes (created using MobileMaker) led to qualitatively different tester feedback compared to that of traditional AI prototypes,
@@
  • a follow-up reflection activity with two designers, showing how designers can make use of fully functional revised prototypes created by testers in the wild,
@@
  • a discussion of implications for future design and prototyping platforms, advocating for systems that prioritize quick iterations and authentic tester feedback for AI experienced in-the-wild.
@@ II-B In Situ User Testing
After developing a prototype, it is crucial to perform user testing and gather feedback [24, 41]. Traditional UI prototyping tools like Figma [42] allow testers to interact with click-through prototypes [42, 43, 44], but these often include only pre-defined input and output examples [20, 45], and the quality and relevance of these examples may not fully represent the diverse contexts and challenges that end-users may face in the real-world [46]. To circumvent the technical demands of building functional AI prototypes while still engaging end-users, designers often conduct user testing with Wizard of Oz setups [47], where a human simulates AI responses [48, 49, 50]. Although helpful, this often fails to accurately simulate the unique errors and behaviors of actual AI systems [51, 52].
To address these challenges, in this work, we focus on supporting “in situ” user testing of AI prototypes, i.e., enabling interactions with MobileMaker prototypes in their intended real-world environments [14, 13, 53, 54]. Specifically, MobileMaker allows users to capture inputs from their immediate surroundings using their mobile devices and view AI model outputs in that same context. We found that this approach not only facilitated the discovery of serendipitous edge cases from real-world inputs (echoing findings from [55, 56] in the context of in situ ML model testing), but also enabled testers to evaluate the AI’s performance more critically, leveraging the rich contextual cues around them.
@@ II-C Iterative Design Based on User Feedback
In HCI, iterative design is recognized as a critical method that emphasizes the repeated cycle of designing, testing, and refining products based on user feedback [57, 58, 59], which is crucial in helping designers identify user pain points and areas for improvements [58, 60]. The length of each iteration in the design process can vary significantly–ranging from a few days in rapid prototyping scenarios to several weeks or months in more sophisticated systems–based on factors like the iteration’s specific objectives, the level of fidelity of the prototype involved, and the available resources [53, 61, 62, 63]. In this work, we explore the opportunity to accelerate these iteration cycles even further by having MobileMaker instantly implement users’ feedback on the spot, thereby enabling them to immediately interact with a revised version of the prototype. This allowed our study participants to more critically reflect on their own feedback and increased their perceived engagement in the design process. Moreover, testers can submit these updated prototypes as “revisions” back to designers, providing a more direct means of conveying user needs in specific contexts.
@@ III Formative Studies & Design Goals
We began by conducting need-finding interviews with two professional UX designers (D1, D2) from a major technology company, both experienced in designing AI features. We explored their workflows for designing, prototyping, and testing LLM-powered features, as well as the challenges they faced. These discussions inspired the idea of rapidly creating LLM-powered mobile prototypes that can be immediately tested and revised in the wild, leading to the initial development of MobileMaker. Throughout the development process, we continued to solicit feedback from the original two designers as well as two additional designers (D3, D4), incorporating their insights and suggestions for improvement.
@@
Based on the data from these formative studies as well as prior work, we identified three design goals for MobileMaker:
@@
  • DG1: Rapid and frictionless authoring of medium-fidelity, mobile AI prototypes. Designers typically prototype AI functionality and UI separately and with different tools. Integrating the AI and UI into a working prototype usually requires significant engineering resources and could often “take weeks if not months” (D2) with considerable coordination and communication overhead. While using early versions of MobileMaker, designers were excited about the possibility of being able to quickly wrap LLM-powered prompts “in an app shell” (D3) to create medium-fidelity prototypes. They noted that this would substantially accelerate prototyping velocity and reduce their reliance on engineers, as their goal was not to create high-fidelity, pixel-perfect prototypes or fully-fledged mobile applications. Three of them were also very interested in the ability to transform ideas into functional prototypes just by describing their ideas in NL.
  • DG2: Lowering the barriers to on-device testing experience and early-stage feedback. Designers typically solicit early-stage feedback on AI prototypes by either showing testers static UI mockups, or by showing example model inputs and outputs via prompt editors. When presented with an early version of MobileMaker, designers found it compelling to potentially “get the best of both worlds” (D4)– a more dynamic and realistic approximation of the AI compared to UI mockups, and a more user-accessible UI compared to the raw model inputs and outputs seen in prompt editors. Furthermore, designers found it compelling to be able to go from an initial idea to a testable mobile interface quickly, as typical mobile development processes can be painstaking and slow. They were excited to have teammates and other users interact with their prototypes on-the-go, rather than being restricted to desktop environments.
@@
  • DG3: Expedited design iteration loop. Traditionally, refining a mobile prototype after user testing is a lengthy and resource-intensive process: designers must make revisions, collaborate with engineers to update the prototype, and schedule new testers, all of which can take days or weeks. To streamline this, designers envisioned a future where the feedback loop is drastically shortened. Designers were excited about the possibility of receiving tester feedback, then having an instant mechanism to produce an improved version of the prototype for another round of testing. This approach could not only keep testers more engaged and motivated compared to traditional methods, but also reduce designers’ dependence on engineers for prototype updates, which typically could take “days or even weeks” (D1).
@@ IV MobileMaker
Based on the above design rationales, we describe the system and user interface design of MobileMaker. Our formative studies revealed that designers sought to make prototypes capable of unveiling user intentions, inputs, and reflections on model outputs without being overly complex. Thus, our objective was not to develop a platform for authoring fully-functional mobile applications; instead, we focused on implementing a minimal set of features essential for designers to explore the human-AI interactions of the prototypes.
@@ IV-A Creating Prototypes
In MobileMaker, designers can build prototypes using three types of configurable UI widgets (Fig. 2):
  • Input Widgets represent different modalities of input that designers can provide, including text, dropdown menus (with a list of predefined text values), live-camera photos (captured live from the front or back camera feed on the mobile device), and file uploads (e.g., photos from an on-device album) (Fig. 2A)
@@
  • Output Widgets are used to display the output from generative models, i.e., LLMs, multimodal LLMs, and image generation models (e.g., [64, 65]) (Fig. 2C).
@@
Each widget type offers customizable visual elements, such as text input placeholders (e.g., “Enter a color”), widget labels (e.g., “Upload a photo of your room”), and the choice to display model outputs as plain text or in a visual component (e.g., carousel card). Designers can also configure various global parameters, include the prototype’s name, description, font style, and layout style, i.e., whether camera controls should be laid out in the list of input controls or as a full screen layout behind all other controls.
@@ IV-B Prototyping with Natural Language
From early feedback, we discovered that designers appreciate the capability to quickly convert ideas into functional mobile prototypes (DG1)–despite MobileMaker’s straightforward interface, some designers found translating high-level concepts to specific low-level widgets cognitively demanding, and could hinder “spur of the moment” ideas (D3). To reduce this barrier, MobileMaker additionally supports prototyping through two NL features:
IV-B1 Creating with Natural Language
In order to allow designers to quickly bootstrap a prototype from NL, they can simply provide a brief sketch of their prototype idea (e.g., “a feature that generates a music playlist from a photo the user takes”). The system then uses a few-shot LLM prompt to instantly generate a prototype. Each example in the few-shot prompt consists of an input NL request and output JSON specification, which encompasses configurations for input, action, and output widgets and as detailed previously. Critically, output widget prompts are dynamically generated to reference the corresponding generated input widgets, reducing the need for further manual configuration.
IV-B2 Revising with Natural Language
To allow testers to rapidly modify prototypes (e.g., in response to in-the-wild inspirations), MobileMaker also supports NL revisions (Fig. 2E). To achieve this, they can simply provide a brief description of the desired changes in NL, such as “add a dropdown for music genre.” The system then categorizes the request as either an update to an existing output widget prompt or an adjustment to the prototype structure (e.g., adding or removing widgets). For the former, the system executes a prompt-revision meta-prompt (see appendix A3) designed to modify an existing prompt based on the NL request. For the latter, the system combines the NL request and the current prototype JSON in a few-shot prompt (similar to the one for NL prototype creation, where each few-shot example includes a revision request, an initial JSON representation, and an updated JSON representation), which outputs a revised JSON specification for the entire prototype.
When someone requests an NL revision, they will receive a summary outlining the changes between the original prototype and the revised versions. They can toggle between these two versions and decide whether to accept the revision. This feature makes it easy for users to quickly compare and confirm if their requested revision has indeed been implemented (e.g., during in-the-wild testing), enhancing their ability to confidently and efficiently iterate on their design.
Together, these NL features enable rapid development and require minimal manual configuration, making MobileMaker’s medium-fi prototyping process accessible and approachable to not only designers but also testers with limited prototyping or prompting experience.